OpenStudy (lcnettech):

please help .will give medal . Sue counted the turtles and birds in the pond. She counted 14 heads and 44 leg. How many turtles did Sue count? A. 4 B. 8 c. 11 D. 22

OpenStudy (gorv):

birds and turtle each have one head

OpenStudy (gorv):

birds have 2 leg and turtle 4

OpenStudy (gorv):

let we have x birds and y turtle so head= x+y=14 legs = 2*x +4*y=44

OpenStudy (lcnettech):

so we cross out d but then we still have 11 8 and 4

OpenStudy (gorv):

x+y=14.........(1) 2x+4y=44 dividing it by 2 x+2y=22........(2) now subtract 1 from 2 x+2y=22 x+y=14 - - - ------------ 0+y=8 so y=8 that is 8 turtle x+8=14 x=6 birds

OpenStudy (lcnettech):

thanks